# TT-BLIP: Enhancing Fake News Detection Using BLIP and Tri-Transformer thanks:

### v1

@@ III-C Feature Fusion Layer
where $\mathbf{W}^{Qt}_{num}\in\mathbb{R}^{d_{t}\times d_{q}}$ is the weight matrix of $Q_{t}$ when computing the head of the $num^{th}$ text modality; $\mathbf{W}^{K*}_{num}\in\mathbb{R}^{d_{*}\times d_{k}}$ be the weight matrix of $K_{*}$ when computing the head of the $num^{th}$** modality; and $\mathbf{W}^{V*}_{num}\in\mathbb{R}^{d_{*}\times d_{v}}$ be the weight matrix of $V_{*}$ when computing the head of the $num^{th}$** modality, where *$*\in\{i,t,m\}$.
Following this, all heads of the *** modalities are connected, denoted as $Y_{*}$, as follows:
@@
Where $\mathbf{W}{O_{*}}$ is the weight matrix multiplied after splicing the head of *** modalities and $n$ denotes the number of self-attention heads used. Thus, the text-based image representation $f_{it}$, the text-based image-text representation $f_{mt}$, and text representation $f_{t}$ can be obtained as follows:

### v2

@@ III-C Feature Fusion Layer
where $\mathbf{W}^{Qt}_{num}\in\mathbb{R}^{d_{t}\times d_{q}}$ is the weight matrix of $Q_{t}$ when computing the head of the $num^{th}$ text modality; $\mathbf{W}^{K*}_{num}\in\mathbb{R}^{d_{*}\times d_{k}}$ be the weight matrix of $K_{*}$ when computing the head of the $num^{th}$ modality; and $\mathbf{W}^{V*}_{num}\in\mathbb{R}^{d_{*}\times d_{v}}$ be the weight matrix of $V_{*}$ when computing the head of the $num^{th}$ modality, where $*\in\{i,t,m\}$.
Following this, all heads of the * modalities are connected, denoted as $Y_{*}$, as follows:
@@
Where $\mathbf{W}{O_{*}}$ is the weight matrix multiplied after splicing the head of * modalities and $n$ denotes the number of self-attention heads used. Thus, the text-based image representation $f_{it}$, the text-based image-text representation $f_{mt}$, and text representation $f_{t}$ can be obtained as follows: